# Bernstein-type Inequalities for Markov Chains and Markov Processes: A Simple and Robust Proof

### v3

@@ 2.3. Bernstein-type tail bound
Let $\left\{Z_{k}\right\}_{k\in\mathbb{N}_{+}}$ be a Markov chain on $\Omega$ with an invariant distribution $\mu$. Suppose that the corresponding transition operator $\bm{P}$ admits an IP gap $\eta_{p}>0$. Let $\nu$ be the initial distribution of $\left\{Z_{k}\right\}_{k\in\mathbb{N}_{+}}$, i.e. $Z_{1}\sim\nu$, such that $\nu$ is absolutely continuous with respect to $\mu$ and that $\nu/\mu$ has a finite $p$-moment for some $p\in[1,+\infty]$. Let $f\in\Omega\to\mathbb{R}$ be a function such that $\operatorname{\mathbb{E}}_{\mu}[f]=0$, $|f|\leq M$, and $\operatorname{Var}_{\mu}[f]\leq\sigma^{2}$. Then, for all $n\in\mathbb{N}_{+}$ and all $\delta\geq 0$,
@@
Recall that the classical Bernstein-type inequalities for general Markov chains, e.g. those obtained in [5, 4], take the form
@@
where $C$ is some universal constant, and $\eta_{a}$ is the absolute spectral gap (1.5). In comparison, our Bernstein-type inequality takes the form
@@
One can see that, when $\delta\ll\sigma\ll M$, (2.16) could be sharper than (2.17). In other cases, the two bounds are comparable. Another important difference between the two results is that (2.16) requires $\bm{P}$ admits a nonzero absolute spectral $\eta_{a}$, while (2.17) only needs a nonzero IP gap $\eta_{p}$. In Appendix A we will show that $\eta_{p}\geq\eta_{a}$ for any Markov transition operator $\bm{P}$. Furthermore, when $\bm{P}$ is a finite-state irreducible transition matrix, the IP gap is always nonzero, but the absolute spectral gap may be zero. Therefore, our result (2.17) is more robust than the classical results (2.16).
@@ 3.2. Exponential moment bound
Again, suppose that $f\in\Omega\to\mathbb{R}$ satisfies $\operatorname{\mathbb{E}}_{\mu}[f]=0$, $|f|\leq M$, and $\operatorname{Var}_{\mu}[f]\leq\sigma^{2}$. The goal of this subsection is to bound the mgf of $\int_{0}^{t}f(Z_{s})\,\mathrm{d}{s}$, which is defined as
@@
$\mathcal{L}G_{t}(z)$ $=\lim_{\tau\rightarrow 0+0}\frac{1}{\tau}\left(\bm{P}_{\tau}G_{t}(z)-G_{t}(z)\right)$
$=\frac{\mathrm{d}{\,}}{\mathrm{d}{\tau}}\operatorname{\mathbb{E}}[G_{t}(Z_{\tau})\,|\,Z_{0}=z]\Big{|}_{\tau=0}$
$=\frac{\mathrm{d}{\,}}{\mathrm{d}{\tau}}\operatorname{\mathbb{E}}\left[\mathrm{e}^{F\left(Z_{[\tau,t+\tau]}\right)}\,\big{|}\,Z_{0}=z\right]\Big{|}_{\tau=0}$
$=\operatorname{\mathbb{E}}\left[(f(Z_{t+\tau})-f(Z_{\tau}))\mathrm{e}^{F\left(Z_{[\tau,t+\tau]}\right)}\,\big{|}\,Z_{0}=z\right]\Big{|}_{\tau=0}$
@@
which is the claimed result. ∎
Now we can bound $a(t)$ using Gronwall’s inequality. To be specific, we have the following lemma.
@@
Let $f:\Omega\to\mathbb{R}$ be a function such that $\operatorname{\mathbb{E}}_{\mu}[f]=0$, $|f|\leq M$, and $\operatorname{Var}_{\mu}[f]\leq\sigma^{2}$. Suppose that $01-4M^{2}/\eta_{p}^{2}>$ 0, then for $t\geq 0$,
@@
Note that $a(0)=1$. Taking the derivative of $a(t)$ for $t\geq 0$ gives
@@
$a(t)\leq a(0)\cdot\mathrm{e}^{2\sigma Mt/(c\eta_{p})}=\mathrm{e}^{2\sigma Mt/(c\eta_{p})},\quad t\geq 0.$
@@
We can then bound the mgf of $\int_{0}^{t}f(Z_{s})\,\mathrm{d}{s}$ for sufficiently small $\theta$, and we omit the obvious proof.
@@ 3.3. Bernstein-type tail bound
Let $(Z_{t})_{t\geq 0}$ be a Markov process on $\Omega$ with an invariant distribution $\mu$. Suppose that the corresponding Markov semigroup $(\bm{P}_{t})_{t\geq 0}$ admits an IP gap $\eta>0$. Let $\nu$ be the initial distribution of $(Z_{t})_{t\geq 0}$, i.e. $Z_{0}\sim\nu$, such that $\nu$ is absolutely continuous with respect to $\mu$ and that $\nu/\mu$ has a finite $p$-moment for some $p\in[1,+\infty]$. Let $f\in\Omega\to\mathbb{R}$ be a function such that $\operatorname{\mathbb{E}}_{\mu}[f]=0$, $|f|\leq M$, and $\operatorname{Var}_{\mu}[f]\leq\sigma^{2}$. Then, for all $t\geq 0$ and all $\delta\geq 0$,
@@
given that $c(q\theta)^{2}=1-4q^{2}\theta^{2}M^{2}/\eta_{p}^{2}\geq 0$. In particular, by choosing
@@ Acknowledgements
The authors are supported by the National Key R&D Program of China under the grant 2021YFA1001500.
@@ Appendix A Comparison of different spectral gaps
To the best of our knowledge, most of the existing results about concentration inequalities for nonreversible Markov chains require a nonzero absolute spectral gap of the transition operator. This condition, however, may fail in some applications.
Another quantity appearing in the literature is the symmetric spectral gap (also called the right spectral gap), which can be interpreted as the spectral gap of the additive symmetrization of $\bm{P}$. To be specific, let $\bm{P}^{*}$ be the adjoint of $\bm{P}$ with respect to the $L_{2,\mu}$ inner product, and let $\bm{A}:=(\bm{P}+\bm{P}^{*})/2$ be the additive symmetrization of $\bm{P}$. We say $\bm{P}$ admits a nonzero symmetric spectral gap $\eta_{s}=1-\lambda_{s}$ if
@@
Here, we argue that for any transition operator $\bm{P}$, the relations $\eta_{p}\geq\eta_{s}\geq\eta_{a}$ always hold. As a direct consequence, our Bernstein-type inequality also holds if $\bm{P}$ admits a nonzero absolute spectral gap or a nonzero symmetric spectral gap.
@@
Taking infimum over $h\in L_{2,\mu}^{0}$, we get $\eta_{s}\leq\eta_{p}$. Similarly, since
@@
holds for any $h\in L_{2,\mu}^{0}$, we deduce $\lambda_{s}\leq\lambda_{a}$, and therefore $\eta_{s}\geq\eta_{a}$. ∎
When the number of states $|\Omega|$ is finite, $\lambda_{s}$ is the second largest eigenvalue of the matrix $\bm{A}$. Moreover, if the transition matrix $\bm{P}$ is irreducible, then it is not hard to see that $\bm{A}$ is also an irreducible transition matrix. Therefore, by the Perron-Frobenius theorem, we always have $\lambda_{s}<1$, and thus $\eta_{p}\geq\eta_{s}=1-\lambda_{s}>0$. On the other hand, by Jensen’s inequality,
$\|\bm{P}h\|_{\mu}^{2}=\sum\limits_{u\in\Omega}\mu(u)\left(\sum\limits_{v\in\Omega}p(u,v)h(v)\right)^{2}\leq\sum\limits_{u\in\Omega}\mu(u)\sum\limits_{v\in\Omega}p(u,v)h(v)^{2}=\|h\|_{\mu}^{2}.$ (A.6)
Hence $\lambda_{a}\leq 1$. However, irreducibility does not guarantee that $\eta_{a}>0$. We give a simple example as follows. Consider the transition matrix
@@
One can check that $\bm{P}$ is irreducible and has a unique invariant distribution $\mu=(1/4,1/4,1/4,1/4)$. It is then apparent that this $\bm{P}$ satisfies the detailed balance condition $\bm{D}_{\mu}\bm{P}=\bm{P}\bm{D}_{\mu}$, where $\bm{D}_{\mu}$ is the diagonal matrix with the elements of $\mu$ on the diagonal. By definition, $\lambda_{a}$ is the second largest singular value of the matrix $\bm{D}_{\mu}^{1/2}\bm{P}\bm{D}_{\mu}^{-1/2}=\bm{P}$. Directly calculation shows that the largest two singular values of $\bm{P}$ are both $1$, and therefore the absolute spectral gap of $\bm{P}$ is 0.
@@ Appendix B Potential proof gap in two related works
In [3, 4], the authors rely on the following lemma to prove concentration inequalities for Markov chains with a nonzero right spectral gap (i.e. the symmetric spectral gap in our notation):
@@
Let $\left\{Z_{k}\right\}_{k\geq 1}$ be a Markov chain with invariant measure $\mu$ and right spectral gap $01-\lambda_{s}>$ 0. Then for any bounded function $f$ and any $t\in\mathbb{R}$,
@@
where $w(\bm{A})$ is the numerical radius of $\bm{A}$: $w(\bm{A}):=\operatorname{sup}\limits_{\|x\|=1}\left|\langle\bm{A}x,x\rangle\right|$.
@@
where $\omega_{k}=\mathrm{e}^{2\pi ki/n}$. For any unit vector $x\in\mathcal{H}$, let $x_{j}=\frac{1}{n}\prod_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{n}(\bm{\mathbf{I}}-\omega_{k}\bm{A})x$. Then we have $x=\sum_{j=1}^{n}x_{j}$. We can compute that
@@
$1-\mathrm{e}^{in\theta}\left\langle\bm{A}^{n}x,x\right\rangle=\sum\limits_{j=1}^{n}\|x_{j}\|^{2}\left(1-\omega_{j}\mathrm{e}^{i\theta}\left\langle\frac{\bm{A}x_{j}}{\|x_{j}\|},\frac{x_{j}}{\|x_{j}\|}\right\rangle\right),\quad\text{forany}\,\ \theta\in\mathbb{R}.$ (B.8)
@@
The above proof does not apply to the real case because we cannot guarantee that for any real unit vector $x$, the corresponding $x_{j}$ in the above proof are all real vectors.

### v4

@@ 2.3. Bernstein-type tail bound
Let $\left\{Z_{k}\right\}_{k\in\mathbb{N}_{+}}$ be a Markov chain on $\Omega$ with an invariant distribution $\mu$. Suppose that the corresponding transition operator $\bm{P}$ admits an IP gap $\eta_{p}>0$. Let $\nu$ be the initial distribution of $\left\{Z_{k}\right\}_{k\in\mathbb{N}_{+}}$, i.e. $Z_{1}\sim\nu$, such that $\nu$ is absolutely continuous with respect to $\mu$ and that $\nu/\mu$ has a finite $p$-moment for some $p\in[1,+\infty]$. Let $f\in\Omega\to\mathbb{R}$ be a function such that $\operatorname{\mathbb{E}}_{\mu}[f]=0$, $|f|\leq M$, and $\operatorname{Var}_{\mu}[f]\leq\sigma^{2}$ for some $M,\sigma\in(0,+\infty)$. Then, for all $n\in\mathbb{N}_{+}$ and all $\delta\geq 0$,
@@
Recall that the classical Bernstein-type inequalities for general Markov chains, e.g. those obtained in [8, 6], take the typical form
@@
where $C$ is some universal constant, and $\eta_{a}$ is the absolute spectral gap. Paulin [13] obtained an improvement of (2.17) in the sense that $\eta_{a}$ is replaced by the pseudo spectral $\eta_{ps}$. In comparison, our Bernstein-type inequality takes the form
@@
In practice, the deviation $\delta$ of the empirical mean is typically of order $1/\sqrt{n}$ and thus is much smaller than $\sigma$ when $n$ is large. Besides, when $\sigma\ll M$, Bernstein-type inequalities ((2.17) and (2.18)) are more preferable than Hoeffding-type inequalities (as those in [11, 15, 4] that also use $\eta_{a}$). For parameters in this range, i.e. $\delta\ll\sigma\ll M$, the traditional result (2.17) (with a variance proxy $\sigma^{2}$) could be sharper than our result (2.18) (with a variance proxy $M\sigma$). However, the significance of our new result (2.18) lies in that it only relies on an IP gap $\eta_{p}$ of $\bm{P}$, while (2.17) relies on an absolute spectral $\eta_{a}$ or a pseudo spectral gap $\eta_{ps}$. In Appendix A, we show that $\eta_{p}\geq\eta_{a}$ and $\eta_{p}\geq\eta_{ps}/2$ for any (non-reversible) Markov transition operator $\bm{P}$. Furthermore, when $\bm{P}$ is a finite-state irreducible transition matrix, $\eta_{p}$ is always nonzero, while $\eta_{a}$ or $\eta_{ps}$ can be arbitrarily small compared to $\eta_{p}$ or even zero (see the examples in Appendix A). In this sense, our result (2.18) is more robust than the classical one (2.17).
Nevertheless, it would still be great to improve our result by establishing a Bernstein inequality in the standard form (2.17) but also with the IP gap $\eta_{p}$ in use. Unfortunately, though this difficulty seems to be merely technical, our current method does not seem to be able to overcome it. We hope to dig into this problem in future works.
@@ 3.2. Exponential moment bound
Again, let $f\in\Omega\to\mathbb{R}$ be such that $\operatorname{\mathbb{E}}_{\mu}[f]=0$, $|f|\leq M$, and $\operatorname{Var}_{\mu}[f]\leq\sigma^{2}$. To avoid non-essential discussions, it is also reasonable to assume that $f(Z_{t})$ is weakly right-continuous in $t$: for any $z\in\Omega$, any $s\geq 0$, and any bounded random variable $X=X(\{Z_{t}\}_{t\geq 0})$,
$\lim_{\tau\rightarrow 0+0}\operatorname{\mathbb{E}}\left[f(Z_{s+\tau})X\big|\,Z_{0}=z\right]=\operatorname{\mathbb{E}}\left[f(Z_{s})X\big|\,Z_{0}=z\right].$ (3.2)
Note that this condition only requires mild regularity of $f$ and $\{Z_{t}\}_{t\geq 0}$ and is satisfied in most of the practical applications.
The goal of this subsection is to bound the mgf of $\int_{0}^{t}f(Z_{s})\,\mathrm{d}{s}$, which is defined as
@@
$\mathcal{L}G_{t}(z)$ $=\lim_{\tau\rightarrow 0+0}\frac{\bm{P}_{\tau}G_{t}(z)-G_{t}(z)}{\tau}$
$=\lim_{\tau\rightarrow 0+0}\operatorname{\mathbb{E}}\left[\frac{G_{t}(Z_{\tau})-G_{t}(Z_{0})}{\tau}\,\Big|\,Z_{0}=z\right]$
$=\lim_{\tau\rightarrow 0+0}\operatorname{\mathbb{E}}\left[\frac{\mathrm{e}^{F(Z_{[\tau,t+\tau]})}-\mathrm{e}^{F(Z_{[0,t]})}}{\tau}\,\Big|\,Z_{0}=z\right]$
@@
which is the claimed result. We have used (3.2) for calculating the limit above. ∎
Next, we can bound $a(t)$ using Grönwall’s inequality. To be specific, we have the following estimate.
@@
Let $f:\Omega\to\mathbb{R}$ be a function such that $\operatorname{\mathbb{E}}_{\mu}[f]=0$, $|f|\leq M$, and $\operatorname{Var}_{\mu}[f]\leq\sigma^{2}$. Suppose that $1-4M^{2}/\eta_{p}^{2}>0$, then for any $t\geq 0$,
@@
Note that $a(0)=1$. Taking the derivative of $a(t)$ (by using (3.2)) for $t\geq 0$ gives
@@
$a(t)\leq a(0)\exp\left(\frac{2\sigma M\,t}{c\,\eta_{p}}\right)=\exp\left(\frac{2\sigma M\,t}{c\,\eta_{p}}\right),\quad t\geq 0.$
@@
We can then bound the mgf of $\int_{0}^{t}f(Z_{s})\,\mathrm{d}{s}$ for sufficiently small $\theta$ again by replacing $f$ with $\theta f$ in Proposition 3.2, and we omit the obvious proof.
@@ 3.3. Bernstein-type tail bound
Let $(Z_{t})_{t\geq 0}$ be a Markov process on $\Omega$ with an invariant distribution $\mu$. Suppose that the corresponding Markov semigroup $(\bm{P}_{t})_{t\geq 0}$ admits an IP gap $\eta_{p}>0$. Let $\nu$ be the initial distribution of $(Z_{t})_{t\geq 0}$, i.e. $Z_{0}\sim\nu$, such that $\nu$ is absolutely continuous with respect to $\mu$ and that $\nu/\mu$ has a finite $p$-moment for some $p\in[1,+\infty]$. Let $f\in\Omega\to\mathbb{R}$ be a function such that $\operatorname{\mathbb{E}}_{\mu}[f]=0$, $|f|\leq M$, and $\operatorname{Var}_{\mu}[f]\leq\sigma^{2}$ for some $M,\sigma\in(0,+\infty)$. Then, for all $t\geq 0$ and all $\delta\geq 0$,
@@
given that $c(q\theta)^{2}=1-4q^{2}\theta^{2}M^{2}/\eta_{p}^{2}>0$. In particular, by choosing
@@
which gives
$c(q\theta)=\frac{2\sigma}{\sqrt{4\sigma^{2}+\delta^{2}}},$
@@ Acknowledgements
The authors are supported by the National Key R&D Program of China under the grant 2021YFA1001500. We are very grateful for the constructive comments on this paper from Joel Tropp, Sam Power, and anonymous referees.
@@ Appendix A Comparison of different spectral gaps
To the best of our knowledge, most of the existing results about concentration inequalities for non-reversible Markov chains require a nonzero absolute spectral gap of the transition operator. This condition, however, may fail in some applications. As a generalization of the absolute spectral gap, Paulin [13] introduced the pseudo spectral gap that is defined as
$\eta_{ps}:=\sup_{k\in\mathbb{N}_{+}}\frac{1-\lambda_{a,k}^{2}}{k},$ (A.2)
where
$\lambda_{a,k}:=\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\|\bm{P}^{k}h\|_{\mu}}{\|h\|_{\mu}}.$ (A.3)
Note that by Jensen’s inequality,
$\|\bm{P}h\|_{\mu}^{2}=\sum\limits_{u\in\Omega}\mu(u)\left(\sum\limits_{v\in\Omega}p(u,v)h(v)\right)^{2}\leq\sum\limits_{u\in\Omega}\mu(u)\sum\limits_{v\in\Omega}p(u,v)h(v)^{2}=\|h\|_{\mu}^{2},$ (A.4)
and thus we have $\lambda_{a,k}\leq\lambda_{a}^{k}\leq 1$ for all $k\in\mathbb{N}_{+}$.
Another quantity appearing in the literature is the symmetric spectral gap (also called the right spectral gap in [6, 4]), which is defined to be the spectral gap of the additive symmetrization of $\bm{P}$. To be specific, let $\bm{P}^{*}$ be the adjoint of $\bm{P}$ with respect to the $L_{2,\mu}$ inner product, and let $\bm{A}:=(\bm{P}+\bm{P}^{*})/2$ be the additive symmetrization of $\bm{P}$. We say $\bm{P}$ admits a nonzero symmetric spectral gap $\eta_{s}=1-\lambda_{s}$ if
@@
Here, we argue that for any transition operator $\bm{P}$, the relations $\eta_{p}\geq\eta_{s}\geq\eta_{a}$ and $\eta_{p}\geq\eta_{ps}/2$ always hold. The proof of $\eta_{p}\geq\eta_{ps}/2$ is due to Chatterjee [1], and we still provide the proof below for the reader’s convenience. As a direct consequence, our Bernstein-type inequality also holds if $\bm{P}$ admits a nonzero absolute spectral gap, a nonzero pseudo spectral gap, or a nonzero symmetric spectral gap.
@@
Taking the infimum over all $h\in L_{2,\mu}^{0}$, we get $\eta_{s}\leq\eta_{p}$. Similarly, since
@@
holds for any $h\in L_{2,\mu}^{0}$, we deduce $\lambda_{s}\leq\lambda_{a}$, and therefore $\eta_{s}\geq\eta_{a}$.
To prove $\eta_{p}\geq\eta_{ps}/2$, we first note that for any $k\in\mathbb{N}_{+}$,
$\|(\bm{P}^{k-1}-\bm{P}^{k})h\|_{\mu}=\|\bm{P}^{k-1}(\mathbf{I}-\bm{P})h\|_{\mu}\leq\|(\mathbf{I}-\bm{P})h\|_{\mu}=\|\bm{L}h\|_{\mu}.$
Then, for any $h\in L_{2,\mu}^{0}$ and any $k\in\mathbb{N}_{+}$, we can compute by triangle inequality and telescoping that
$1-\frac{\|\bm{P}^{k}h\|_{\mu}}{\|h\|_{\mu}}\leq\frac{\|(\mathbf{I}-\bm{P}^{k})h\|_{\mu}}{\|h\|_{\mu}}\leq\frac{1}{{\|h\|_{\mu}}}\sum\limits_{i=1}^{k}\|(\bm{P}^{i-1}-\bm{P}^{i})h\|_{\mu}\leq k\frac{\|\bm{L}h\|_{\mu}}{{\|h\|_{\mu}}}.$
It follows by Taking the infimum over all $h\in L_{2,\mu}^{0}$ that $1-\lambda_{a,k}\leq k\eta_{p}$, and thus
$\eta_{ps}=\sup_{k\in\mathbb{N}_{+}}\frac{1-\lambda_{a,k}^{2}}{k}\leq\sup_{k\in\mathbb{N}_{+}}\frac{2(1-\lambda_{a,k})}{k}\leq 2\eta_{p}.$
We have used above the fact that $\lambda_{a,k}\leq 1$ for all $k\in\mathbb{N}_{+}$. ∎
When the number of states $|\Omega|$ is finite, $\lambda_{s}$ is the second largest eigenvalue of the matrix $\bm{A}$. Moreover, if the transition matrix $\bm{P}$ is irreducible, then it is not hard to see that $\bm{A}$ is also an irreducible transition matrix. Therefore, by the Perron–Frobenius theorem, we always have $\lambda_{s}<1$, and thus $\eta_{p}\geq\eta_{s}=1-\lambda_{s}>0$. However, irreducibility does not guarantee a uniform lower bound of the ratios $\eta_{a}/\eta_{p}$ and $\eta_{ps}/\eta_{p}$. Below are some simple examples demonstrating the superiority of the IP gap.
Example A.2.
Chatterjee provided in [1] an example demonstrating that the ratio $\eta_{a}/\eta_{p}$ can be arbitrarily small. Consider the one dimensional random walk on the periodic grid $\mathbb{Z}/n\mathbb{Z}$ where the probability of not moving and that of moving one step to the right are both $1/2$. It is then not hard to check that the IP gap $\eta_{p}$ of this Markov chain is of order $1/n$, while the symmetric spectral gap $\eta_{s}$ and the absolute spectral $\eta_{a}$ are both of order $1/n^{2}$.
Example A.3.
The next example (modified from Chatterjee’s random walk example) also shows the superiority of the IP gap over the other gaps that have been used to establish concentration inequalities for non-reversible Markov chains. Consider a parameterized family of transition matrices,
$\bm{P}_{0}=\begin{bmatrix}0&1&0\\ 0&0&1\\ 1&0&0\\ \end{bmatrix}\quad\text{and}\quad\bm{P}_{\varepsilon}=\varepsilon\mathbf{I}+(1-\varepsilon)\bm{P}_{0},\quad\varepsilon\in[0,1].$ (A.8)
Suppose $\varepsilon\ll 1$. One can easily show by a straightforward calculation and the spectral perturbation theory that
$\eta_{p}(\bm{P}_{\varepsilon})=(1-\varepsilon)\sqrt{3},\quad\eta_{a}(\bm{P}_{\varepsilon})=O(\varepsilon),\quad\eta_{ps}(\bm{P}_{\varepsilon})=O(\varepsilon).$
In particular, $\eta_{p}(\bm{P}_{0})=\sqrt{3}$ and $\eta_{a}(\bm{P}_{0})=\eta_{ps}(\bm{P}_{0})=0$.
Example A.4.
Here, we give another simple example where $\eta_{p}>0$ but $\eta_{a}=0$. Consider the transition matrix
@@
One can check that $\bm{P}$ is irreducible and has a unique invariant distribution $\mu=(1/4,1/4,1/4,1/4)$. By definition, $\lambda_{a}$ is the second largest singular value of the matrix $\bm{D}_{\mu}^{1/2}\bm{P}\bm{D}_{\mu}^{-1/2}$, where $\bm{D}_{\mu}$ is the diagonal matrix with the elements of $\mu$ on the diagonal. Since $\bm{D}_{\mu}$ is a multiple of the identity matrix, $\bm{D}_{\mu}^{1/2}\bm{P}\bm{D}_{\mu}^{-1/2}=\bm{P}$. A direct calculation shows that the largest two singular values of $\bm{P}$ are both $1$, and therefore the absolute spectral gap of $\bm{P}$ is 0.
@@ Appendix B Potential proof gap in two related works
In [4, 6], the authors rely on the following lemma to establish concentration inequalities for non-reversible Markov chains with a nonzero symmetric spectral gap $\eta_{s}$ (i.e. the right spectral gap in their notation):
@@
Let $\left\{Z_{k}\right\}_{k\geq 1}$ be a Markov chain with invariant measure $\mu$ and right spectral gap $1-\lambda_{s}>0$. Then for any bounded function $f$ and any $t\in\mathbb{R}$,
@@
where $w(\bm{A})$ is the numerical radius of $\bm{A}$: $w(\bm{A}):=\sup_{\|x\|=1}\left|\langle\bm{A}x,x\rangle\right|$.
@@
where $\omega_{k}=\mathrm{e}^{2\pi k\mathrm{i}/n}$. For any unit vector $x\in\mathcal{H}$, let
$x_{j}=\frac{1}{n}\prod\limits_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{n}(\bm{\mathbf{I}}-\omega_{k}\bm{A})x.$
Then we have $x=\sum_{j=1}^{n}x_{j}$. We can compute that
@@
[…]
@@
To prove a real version of Theorem B.2, one must only assume and use that $\left|\langle\bm{A}x,x\rangle\right|\leq 1$ for all real $x$ with $\|x\|=1$. However, for $n\geq 3$, the polynomial $1-z^{n}$ always has complex roots, and thus the $x_{j}$ defined in the proof above cannot always be real in general. Therefore, one is not able to control $\langle\bm{A}x_{j},x_{j}\rangle$ on the right-hand side in (B.8).